### v2

# Copresheaf Topological Neural Networks: A Generalized Deep Learning Framework

Abstract
We introduce copresheaf topological neural networks (CTNNs), a powerful and unifying framework that encapsulates a wide spectrum of deep learning architectures, designed to operate on structured data: including images, point clouds, graphs, meshes, and topological manifolds. While deep learning has profoundly impacted domains ranging from digital assistants to autonomous systems, the principled design of neural architectures tailored to specific tasks and data types remains one of the field’s most persistent open challenges. CTNNs address this gap by grounding model design in the language of copresheaves, a concept from algebraic topology that generalizes and subsumes most practical deep learning models in use today. This abstract yet constructive formulation yields a rich design space from which theoretically sound and practically effective solutions can be derived to tackle core challenges in representation learning: long-range dependencies, oversmoothing, heterophily, and non-Euclidean domains. Our empirical results on structured data benchmarks demonstrate that CTNNs consistently outperform conventional baselines, particularly in tasks requiring hierarchical or localized sensitivity. These results underscore CTNNs as a principled, multi-scale foundation for the next generation of deep learning architectures.
1 Introduction
Deep learning has excelled by exploiting structural biases—convolutions for images (Krizhevsky et al., 2012), transformers for sequences (Vaswani et al., 2017), and message-passing for graphs (Gilmer et al., 2017). Yet, the design of architectures that generalize across domains with complex, irregular, or multiscale structure remains a notorious challenge (Bronstein et al., 2017; Hajij et al., 2023b). Real-world data—spanning physical systems, biomedical signals, and scientific simulations—rarely adhere to the regularity assumptions embedded in conventional architectures. These data are inherently heterogeneous, directional, and hierarchical, often involving relations beyond pairwise connections or symmetric neighborhoods.
Convolutional networks, designed for uniform grids, gloss over local irregularities; graph neural networks (GNNs) often rely on homophily and oversmooth features as depth increases; and transformers, while excellent at capturing long-range dependencies, assume homogeneous embedding spaces, incur quadratic complexity, and lack built-in notions of anisotropy or variable local structures. These shortcomings underscore the need for a framework that can natively encode diverse local behaviors, respect directional couplings, and propagate information across scales without sacrificing critical distinctions or imposing artificial uniformity.
To address this foundational gap, we propose copresheaf topological neural networks (CTNNs), a unifying framework for deep learning based on copresheaves—a categorical structure that equips each local region of a domain with its own feature space, along with learnable maps specifying how information flows across regions. Unlike traditional models that assume a global latent space and isotropic propagation, our framework respects local variability in representation and directional flow of information, enabling architectures that are inherently multiscale, anisotropic, and expressive.
Figure 1: A copresheaf topological neural network (CTNN) operates on combinatorial complexes (CCs), which generalize Euclidean grids, graphs, meshes, and hypergraphs. A CTNN is characterized by a set of locally indexed copresheaf maps $\rho_{x_{i}\to x_{j}}$, defined between cells $x_{i}$ and $x_{j}$ in the CC, and directed from $x_{i}$ to $x_{j}$. The figure illustrates how a CTNN updates a local representation $\mathbf{h}_{x}$ of a cell $x$ using neighborhood representations $\mathbf{h}_{y}$ and $\mathbf{h}_{z}$, which are sent to $x$ via the learnable local copresheaf maps $\rho_{x\to x}$, $\rho_{y\to x}$ and $\rho_{z\to x}$.
By constructing our CTNNs upon combinatorial complexes (Hajij et al., 2023b, a), which generalize graphs, simplicial complexes, and cell complexes, we enable a principled message-passing mechanism over general topological domains grounded in the theory of copresheaves. This unified perspective subsumes and extends many deep learning paradigms—including GNNs, attention mechanisms, sheaf neural networks (Hansen and Ghrist, 2019d; Bodnar et al., 2022a), and topological neural networks (Papillon et al., 2023a; Hajij et al., 2023b; Bodnar et al., 2021; Ebli et al., 2020; Giusti et al., 2022b) within a single formalism. Our approach further departs from the traditional assumption of a single shared latent space by modeling task-specific, directional latent spaces that bridge diverse deep learning frameworks. Furthermore, CTNNs flexibly handle both Euclidean and non-Euclidean data, supporting expressive architectures such as Copresheaf GNNs, Transformers, and Convolutional Networks, which learn structure-aware, directional transport maps. Overall, CTNNs offer a promising framework for developing next-generation, topologically grounded, structure-aware machine learning models. See Figure 1 for an illustration.
Related work
Our work here is closely related to sheaf neural networks (SNNs) extend traditional graph neural networks by leveraging the mathematical framework of cellular sheaves to capture higher-order and heterogeneous relationships. Early work by Hansen and Ghrist (2019a, b) introduced methods to learn sheaf Laplacians from smooth signals and developed a spectral theory that connects sheaf topology with graph structure. Building on these ideas, Hansen and Gebhart (2020a) proposed the first SNN architecture, demonstrating that incorporating edge-specific linear maps can improve performance on tasks involving asymmetric and heterogeneous relations.
Recent advances have focused on mitigating common issues in graph neural networks such as oversmoothing and heterophily. For instance, Bodnar et al. (2022c) introduced neural sheaf diffusion processes that address these challenges by embedding topological constraints into the learning process. Similarly, Barbero et al. (2022a, b) developed connection Laplacian methods and attention-based mechanisms that further enhance the expressiveness and efficiency of SNNs. The versatility of the sheaf framework has also been demonstrated through its extension to hypergraphs and heterogeneous graphs (Duta et al., 2023; Braithwaite et al., 2024), which enables modeling of complex, higher-order interactions. Moreover, novel approaches incorporating joint diffusion processes (Caralt et al., 2024) and Bayesian formulations (Gillespie et al., 2024) have improved the robustness and uncertainty quantification of SNNs. Finally, the application of sheaf neural networks in recommender systems (Purificato et al., 2024) exemplifies their practical utility in real-world domains. Together, these contributions underscore the potential of SNNs to enrich graph-based learning by integrating topological and geometric information directly into neural architectures. Our proposed CTNNs subsume these architectures while avoiding restrictive co-boundary maps or rank-specific Laplacian operators. The appendix provides a more thorough literature review of related work.
2 Preliminaries
This section presents preliminary concepts needed for developing our theoretical framework. It recalls combinatorial complexes and neighborhood structures, reviews sheaves and copresheaves on directed graphs, and contrasts cellular sheaves with copresheaves in graph-based modeling.
2.1 Combinatorial Complexes and Neighborhood Structures
To ensure utmost generality, we base our framework on the combinatorial complexes (CCs) (Hajij et al., 2023b, a), which unify set-type and hierarchical relations over which data are defined. CC-neighborhood functions then formalize local interactions forming an ideal ground to define sheaves and higher-order message passing schemes for our copresheaf topological neural networks (CTNNs).
Definition 1 (Combinatorial complex (Hajij et al., 2023b)).
A combinatorial complex (CC) is a triple $(\mathcal{S},\mathcal{X},\operatorname{rk})$, where $\mathcal{S}$ is a finite non-empty set of vertices, $\mathcal{X}\subset\mathcal{P}(\mathcal{S})\setminus\{\emptyset\}$, with $\mathcal{P}(\mathcal{S})$ denoting the power set of $\mathcal{S}$, $\operatorname{rk}:\mathcal{X}\to\mathbb{Z}_{\geq 0}$ is a rank function such that $\{s\}\in\mathcal{X}$, $\operatorname{rk}(\{s\})=0$ for all $s\in\mathcal{S}$, and $x\subseteq y\implies\operatorname{rk}(x)\leq\operatorname{rk}(y)$ for all $x,y\in\mathcal{X}$.
Figure 2: A combinatorial complex of dimension 2.
When context permits, we write a CC $(\mathcal{S},\mathcal{X},\operatorname{rk})$ simply as $\mathcal{X}$; each $x\in\mathcal{X}$ has rank $\operatorname{rk}(x)$, $\dim\mathcal{X}=\max_{x\in\mathcal{X}}\operatorname{rk}(x)$, its $k$-cells $x^{k}$ are defined to be the cells $x$ with $\operatorname{rk}(x)=k$, furthermore we set $\mathcal{X}^{k}=\{x\in\mathcal{X}:\operatorname{rk}(x)=k\}=\operatorname{rk}^{-1}(\{k\})$. See Fig 2 for an example.
Definition 2 (Neighborhood function).
Given a CC $\mathcal{X}$, a neighborhood function on a CC $(\mathcal{S},\mathcal{X},\operatorname{rk})$ is a map $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$, which assigns to each cell $x$ in $\mathcal{X}$ a collection of neighbor cells $\mathcal{N}(x)\subset\mathcal{X}$, referred to as the neighborhood of $x$. In our context, two neighborhood functions are common, namely the adjacency neighborhood and the incidence neighborhood: (i) the adjacency neighborhood $\mathcal{N}_{\mathrm{adj}}(x)=\{\,y\in\mathcal{X}\mid\exists\,z\in\mathcal{X}:x\subset z,\;y\subset z\}$ and (ii) the incidence neighborhood $\mathcal{N}_{\mathrm{inc}}(x)=\{\,y\in\mathcal{X}\mid x\subset y\}$, potentially defined with respect to particular ranks.
In practice, neighborhood functions are stored via matrices called neighborhood matrices.
Definition 3 (Neighborhood matrix).
Let $\mathcal{N}$ be a neighborhood function on a CC $\mathcal{X}$. Let $\mathcal{Y}=\{y_{1},\dots,y_{n}\}\subset\mathcal{\mathcal{X}}$, $\mathcal{Z}=\{z_{1},\dots,z_{m}\}\subset\mathcal{\mathcal{X}}$ be two collections of cells s.t. $\mathcal{N}(y_{j})\subseteq\mathcal{Z}$ for all $1\leq j\leq n$. An element of the neighborhood matrix $\mathbf{G}\in\{0,1\}^{m\times n}$ is defined as $[\mathbf{G}]_{ij}=\mathbbm{1}_{\mathcal{N}(y_{j})}\left(z_{i}\right)$.
The copresheaf structure that we develop on CCs depends on the neighborhood function. To introduce it, we first review sheaves and copresheaves on graphs, and then extend these notions to CCs.
2.2 Sheaves and Copresheaves on Directed Graphs
The copresheaf formalism assigns each vertex its unique, natural feature space $\mathcal{F}(x)$, respecting the potentially heterogeneous nature of the data, and each directed edge a transformation $\rho_{x\!\to\!y}$ that tells how data move between those spaces, $\mathcal{F}(x)\to\mathcal{F}(y)$. This separation of where data reside from how they travel provides a principled backbone for learning beyond the single-latent-space assumption of standard deep learning architectures. Concretely, copresheaves are defined as follows [^1] While we avoid overly complicated jargon, the appendix links our constructs to those of category theory for a more rigorous exposition.:
Definition 4 (Copresheaf).
A copresheaf $(\mathcal{F},\rho,G)$ on a directed graph $G=(V,E)$ is given by
  • a real vector space $\mathcal{F}(x)$ for every vertex $x\in V$;
  • a linear map $\rho_{x\!\to\!y}:\mathcal{F}(x)\longrightarrow\mathcal{F}(y)$ for every directed edge $x\!\to\!y\in E$.
Think of a copresheaf as a system for sending messages across a network, where each node has its own language (stalk), and edges translate messages (linear maps) to match the recipient’s language. More concretely, on a directed graph $G=(V,E)$, each vertex $x\in V$ carries a task-specific latent space $\mathcal{F}(x)$, and every edge $x\!\to\!y\in E$ applies a learnable, edge-indexed linear map $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ that re-embeds $x$’s features into $y$’s coordinate frame, thereby realizing directional, embedding-level message passing throughout the network.
Cellular sheaves and copresheaves: contrasting graph learning
While much of the recent literature has focused on sheaf learning (Ayzenberg et al., 2025), our approach is rooted in the copresheaf perspective. Copresheaves offer a novel perspective on deep learning by assigning each vertex $x$ a feature space $\mathcal{F}(x)$, and each edge $x\to y$ a linear map $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ that governs data propagation. This setup departs from the traditional deep learning’s core assumption of a single, shared latent space, enabling the modeling of heterogeneous, task-specific latent spaces and directional relations. The significance of this approach lies in its ability to generalize and connect different deep learning paradigms. Copresheaf-type architectures extend beyond sheaf neural networks (Hansen and Gebhart, 2020b; Bodnar et al., 2022b; Barbero et al., 2022c; Duta et al., 2024; Battiloro et al., 2024a) and topological neural network architectures (Papillon et al., 2023b)—typically designed for non-Euclidean data—by also accommodating Euclidean data effectively. This versatility allows them to unify applications across diverse data domains and architectural frameworks, providing a cohesive structure that leverages directional information flow and adapts to task-specific requirements.
In graph-based modeling, cellular sheaves provide a formal framework to ensure data consistency in undirected graphs by encoding symmetric local-to-global relations via an incidence structure between vertices and edges. These structural points, formalized in the following definition, are encoded through restriction maps, ensuring data consistency between vertices and incident edges.
Definition 5 (Cellular sheaf).
Let $G=(V,E)$ be a undirected graph. Let $x\trianglelefteq e$ indicate that vertex $x\in V$ is incident to edge $e\in E$. A cellular sheaf on $G$ consists of:
  • a vector space $\mathcal{F}(x)$ to each vertex $x\in V$;
  • a vector space $\mathcal{F}(e)$ to each edge $e\in E$;
  • a linear restriction map $\mathcal{F}_{x\trianglelefteq e}:\mathcal{F}(x)\to\mathcal{F}(e)$ for each incidence $x\trianglelefteq e$.
The data on nodes $x$ and $y$, denoted by $\mathbf{h}_{x}\in\mathcal{F}(x)$ and $\mathbf{h}_{y}\in\mathcal{F}(y)$, “agree” on the edge $e$ if their images under the restriction maps coincide:
$\mathcal{F}(x)\xrightarrow{\mathcal{F}_{x\triangleleft e}}\mathcal{F}(e)\xleftarrow{\mathcal{F}_{y\triangleleft e}}\mathcal{F}(y).$ (1)
A global section of a sheaf on a graph $G$ assigns data $\mathbf{h}_{v}\in\mathcal{F}(v)$ to each vertex $v$ and $\mathbf{h}_{e}\in\mathcal{F}(e)$ to each edge $e$, such that for every edge $e$ between nodes $x$ and $y$: $\mathcal{F}_{x\triangleleft e}(\mathbf{h}_{x})=\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})$. This consistency condition ensures data consistency across local connections in a network. Global sections represent equilibrium states where this local agreement holds across the entire graph, enabling unified and coherent data representations for the consensus of a complex system. Most sheaf-based architectures have focused on diffusion-type models (Hansen and Gebhart, 2020b; Bodnar et al., 2022b; Barbero et al., 2022c; Duta et al., 2024; Battiloro et al., 2024a), where the sheaf Laplacian $\Delta_{\mathcal{F}}$ minimizes the Dirichlet energy, thereby enforcing global consistency condition. More precisely, for some arbitrary choice of orientation for each edge, define the coboundary map
$\delta:C^{0}\to C^{1},\qquad(\delta\mathbf{h})_{e}=\mathcal{F}_{x\triangleleft e}(\mathbf{h}_{x})\;-\;\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})\quad\text{for }e=(x,y),$ (2)
which measures local disagreement with respect to the edge $e$. The sheaf Laplacian $\Delta_{\mathcal{F}}=\delta^{T}\delta$ aggregates all the restriction maps $\{\mathcal{F}_{x\triangleleft e}\}$ into a single symmetric, positive semidefinite operator. Its associated quadratic form, $\mathbf{h}^{T}\Delta_{\mathcal{F}}\,\mathbf{h}$, has a trace that defines the sheaf Dirichlet energy.
Unlike sheaves, which enforce data consistency across overlaps, copresheaves model directional data flow—ideal for processes like information propagation, causality, and hierarchical dependencies. They assign vector spaces $F(x)$ only to vertices and define learnable linear maps $\rho_{x\to y}\colon\mathcal{F}(x)\to\mathcal{F}(y)$ along directed edges, without imposing sheaf consistency conditions. This vertex-centric, anisotropic framework naturally integrates with message-passing methods such as GNNs and TNNs, allowing parameterized maps to adapt during training.
3 Copresheaf Topological Neural Networks
We are now ready to introduce our copresheaf topological neural networks (CTNNs), a higher-order message-passing mechanism that generalizes the modeling of relational structures within TNNs as illustrated in Figure 1. We begin by exposing copresheaves induced by neighborhood functions on CCs, providing a structured way to model general, local-to-global relations.
Let $\mathcal{X}$ be a CC and let $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ be a neighborhood function. We define the effective support of $\mathcal{N}$ as the set $\mathcal{X}_{\mathcal{N}}:=\{x\in\mathcal{X}\mid\mathcal{N}(x)\neq\emptyset\}.$ This set identifies cells that receive input from neighbors. The neighborhood function $\mathcal{N}$ induces a directed graph $G_{\mathcal{N}}=(V_{\mathcal{N}},E_{\mathcal{N}})$, where the vertex set is:
$V_{\mathcal{N}}:=\mathcal{X}_{\mathcal{N}}\cup\bigcup_{x\in\mathcal{X}}\mathcal{N}(x),$
and the edge set is:$E_{\mathcal{N}}:=\{y\to x\mid x\in\mathcal{X},y\in\mathcal{N}(x)\}.$ The vertex set $V_{\mathcal{N}}$ includes both cells with non-empty neighborhoods (targets) and their neighbors (sources). This graph determines how data propagates across the complex, with each edge encoding a directional relation from neighbor $y$ to target $x$.
Definition 6 (Neighborhood-dependent copresheaf).
Let $\mathcal{X}$ be a CC, $\mathcal{N}$ a neighborhood function, and $G_{\mathcal{N}}=(V_{\mathcal{N}},E_{\mathcal{N}})$ the induced directed graph with $V_{\mathcal{N}}=\mathcal{X}_{\mathcal{N}}\cup\bigcup_{x\in\mathcal{X}}\mathcal{N}(x)$. An $\mathcal{N}$-dependent copresheaf assigns a vector space $\mathcal{F}(x)$ to each $x\in V_{\mathcal{N}}$, and a linear map $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ for each edge $y\to x\in E_{\mathcal{N}}$.
When clear from context, we simplify the notation from $\mathcal{F}^{\mathcal{N}}$ to $\mathcal{F}$ and $\rho_{y\to x}^{\mathcal{N}}$ to $\rho_{y\to x}$.
Copresheaf neighborhood matrices
Having introduced neighborhood matrices as binary encodings of local interactions, we now generalize this notion to define copresheaf neighborhood matrices (CNMs). Instead of binary entries, CNMs consist of copresheaf maps between data assigned to the cells in a CC, allowing richer encoding of local dependencies. Subsequently, we define specialized versions, such as copresheaf adjacency and incidence matrices, that capture specific topological relations, facilitating structured message-passing in our CTNNs.
In particular, let us define the $k$-cochain space to be the direct sum $C^{k}(\mathcal{X},\mathcal{F}^{\mathcal{N}})=\bigoplus_{x\in\mathcal{X}^{k}}\mathcal{F}^{\mathcal{N}}(x)$, and denote by $\mathrm{Hom}(\mathcal{F}^{\mathcal{N}}(i),\mathcal{F}^{\mathcal{N}}(j))$ the space of maps from the data at $i$ ($i$-th stalk) to that at $j$ ($j$-th stalk). Here, the maps encode how data is transferred or transformed between neighboring cells. We next define the CNM.
Definition 7 (Copresheaf neighborhood matrices).
For a CC $\mathcal{X}$ equipped with a neighborhood function $\mathcal{N}$, let $\mathcal{F}\colon G_{\mathcal{N}}\to\mathbf{Vect}_{\mathbb{R}}$ be a copresheaf on $\mathcal{X}$. Also let $\mathcal{Y}=\{y_{1},\dots,y_{n}\}\subseteq\mathcal{X}$ and $\mathcal{Z}=\{z_{1},\dots,z_{m}\}\subseteq\mathcal{X}$ be two collections of cells s. t. $\mathcal{N}(y_{j})\subseteq\mathcal{Z}$ for all $1\leq j\leq n$. The copresheaf neighborhood matrix of $\mathcal{N}$ with respect to $\mathcal{Y}$ and $\mathcal{Z}$ is the $m\times n$ matrix $\mathbf{G}^{\mathcal{N}}$:
$[\mathbf{G}^{\mathcal{N}}]_{ij}\;=\;\begin{cases}\rho_{\,z_{i}\to y_{j}}\;\in\;\mathrm{Hom}\bigl{(}\mathcal{F}(z_{i}),\mathcal{F}(y_{j})\bigr{)},&z_{i}\in\mathcal{N}(y_{j}),\\ 0,&\text{otherwise}.\end{cases}$ (3)
The neighborhood function $\mathcal{N}$ determines the directed relationships between the cells, and the CNM encodes the interactions between cells induced by these relationships, collecting the maps of a cell from its neighbors. Next, analogous to Dfn. 2, we define copresheaf adjacency and copresheaf incidence matrices as specialized forms of the general CNM.
Definition 8 (Copresheaf adjacency/incidence matrices).
Fix $r,k$. Define $\mathcal{N}_{\mathrm{adj}}^{(r,k)}(x)=\{y\in\mathcal{X}^{r}\mid\exists z\in\mathcal{X}^{r+k},x\preceq z,y\preceq z\}$ and $\mathcal{N}_{\mathrm{inc}}^{(r,k)}(x)=\{y\in\mathcal{X}^{k}\mid x\preceq y\}$. Then the copresheaf adjacency matrix (CAM) $\mathbf{A}_{r,k}\in\mathbb{R}^{|\mathcal{X}^{r}|\times|\mathcal{X}^{r}|}$ and copresheaf incidence matrix (CIM) $\mathbf{B}_{r,k}\in\mathbb{R}^{|\mathcal{X}^{k}|\times|\mathcal{X}^{r}|}$ are
$[\mathbf{A}_{r,k}]_{ij}$ =$=\begin{cases}\rho_{y_{i}\to x_{j}}\in\mathrm{Hom}(\mathcal{F}(y_{i}),\mathcal{F}(x_{j}))&\text{if }y_{i}\in\mathcal{N}_{\mathrm{adj}}^{(r,k)}(x_{j}),\\ 0&\text{otherwise}.\end{cases}$ (4)
$[\mathbf{B}_{r,k}]_{ij}$ =$=\begin{cases}\rho_{z_{i}\to y_{j}}\in\mathrm{Hom}(\mathcal{F}(z_{i}),\mathcal{F}(y_{j}))&\text{if }z_{i}\in\mathcal{N}_{\mathrm{inc}}^{(r,k)}(y_{j}),\\ 0&\text{otherwise}.\end{cases}$ (5)
These specialized matrices capture distinct topological relations. CAM encodes relations in which cells share an upper cell, while the CIM encodes relations where one cell is incident to another.
Copresheaf-based message passing
We now generalize traditional graph message-passing to heterogeneous and higher-order interactions involving cells of varying ranks or multi-way relations going beyond pairwise connections, by explicitly incorporating copresheaf structures defined over CCs. This leads to copresheaf-based message-passing, a flexible and expressive tool for capturing complex, multi-scale relations in structured data.
Proposition 1 (Copresheaf message passing on directed graphs).
Let $G=(V,E)$ be a directed graph and $\mathcal{F},\rho,G$ a copresheaf. For each layer $l$ and vertex $x\in V$, let $\mathbf{h}_{x}^{(l)}\in\mathcal{F}(x)$. Then
$\mathbf{h}_{x}^{(l+1)}=\beta\Bigl{(}\mathbf{h}_{x}^{(l)},\bigoplus_{(y\to x)\in E}\alpha\bigl{(}\mathbf{h}_{x}^{(l)},\rho_{y\to x}\mathbf{h}_{y}^{(l)}\bigr{)}\Bigr{)},$
where $\rho_{y\to x}\colon\mathcal{F}(y)\to\mathcal{F}(x)$ is the linear map associated with edge $y\to x$, $\alpha$ is a learnable message function, $\oplus$ a permutation-invariant aggregator, and $\beta$ a learnable update function.
We refer to to any neural network constructed using Prop. 1 as a copresheaf message-passing neural network (CMPNN). Proposition 1 establishes a unifying framework that generalizes graph-based message passing by incorporating learnable, anisotropic linear maps associated with directed edges. This formulation subsumes many standard graph neural networks (GNNs) such as graph convolutional networks (Kipf and Welling, 2017b), graph attention networks (Veličković et al., 2018), and their variants (Veličković, 2022), by viewing message passing as copresheaf maps on directed graphs. Consequently, all architectures derived from these foundational GNN models naturally fit within our copresheaf message-passing paradigm. Moreover, sheaf neural networks (SNNs), which operate on cellular sheaves over undirected graphs, can be adapted into this framework by reinterpreting their edge-mediated transport operator which as direct vertex-to-vertex morphisms on a bidirected graph. This adaptation is detailed in the following theorem.
Theorem 1.
Let $G=(V,E)$ be an undirected graph equipped with a cellular sheaf $\mathcal{F}$ assigning vector spaces to vertices and edges, and linear maps $\mathcal{F}_{v\trianglelefteq e}$ for each vertex $v\in e$. Then for each edge $e=\{u,v\}\in E$, the SNN message passing from $v$ to $u$, given by the composition $\mathcal{F}_{u\trianglelefteq e}^{\top}\circ\mathcal{F}_{v\trianglelefteq e}:\mathcal{F}(v)\to\mathcal{F}(u),$ can be realized as a single morphism $\rho_{v\to u}:\mathcal{F}(v)\to\mathcal{F}(u)$ in a copresheaf on the bidirected graph $G^{\prime}=(V,E^{\prime}),\quad E^{\prime}=\{(u,v),(v,u)\mid\{u,v\}\in E\},$ by setting $\rho_{v\to u}=\mathcal{F}_{u\trianglelefteq e}^{\top}\circ\mathcal{F}_{v\trianglelefteq e}.$
Theorem 1 demonstrates that the message-passing scheme employed by SNNs can be interpreted as a special case of CMPNNs when restricted to bidirected graphs. This perspective generalizes most existing SNN architectures found in the literature, including those in Hansen and Gebhart (2020b); Bodnar et al. (2022b); Barbero et al. (2022c). The map $\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}$ arises from composing the restriction map $\mathcal{F}_{v\triangleleft e}\colon\mathcal{F}(v)\to\mathcal{F}(e)$ with its adjoint $\mathcal{F}_{u\triangleleft e}^{\top}\colon\mathcal{F}(e)\to\mathcal{F}(u)$, thereby capturing both vertex-to-edge and edge-to-vertex transformations defined by the cellular sheaf. As a consequence of this connection between SNNs and CMPNNs, diffusion-style updates—commonly used in sheaf-based models, such as those based on the sheaf Laplacian—can be succinctly expressed within the CMPNN framework. This result is formally stated in Proposition 2, with a full proof provided in the appendix.
Proposition 2 (Neural-sheaf diffusion (Bodnar et al., 2022b) as copresheaf message passing).
Let $G=(V,E)$ be an undirected graph endowed with a cellular sheaf $\mathcal{F}$. Given vertex features $\mathbf{H}=[\mathbf{h}_{v}]_{v\in V}$ with $\mathbf{h}_{v}\in\mathcal{F}(v)$, and learnable linear maps $W_{1},W_{2}$, define the diffusion update:
$\mathbf{H}^{+}=\mathbf{H}-\bigl{(}\Delta_{\mathcal{F}}\otimes I\bigr{)}(I_{n}\otimes W_{1})\mathbf{H}W_{2},$ (6)
where $\Delta_{\mathcal{F}}=[L_{F,v,u}]_{v,u\in V}$ has blocks $L_{F,vv}=\sum_{v\trianglelefteq e}\mathcal{F}_{v\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$, $L_{F,vu}=-\mathcal{F}_{v\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},\,\text{for}~{}u\neq v,\,u\trianglelefteq e,\,v\trianglelefteq e$. Then, $\mathbf{H}^{+}$ can be expressed in the copresheaf message-passing form of Prp. 1.
The next definition formalizes the notion of general multi-way propagation.
Proposition 3 (Copresheaf-based higher-order message passing).
Let $\mathcal{X}$ be a CC, and $\mathfrak{N}=\{\mathcal{N}_{k}\}_{k=1}^{n}$ a collection of neighborhood functions. For each $k$, let $(\mathcal{F}^{\mathcal{N}_{k}},\rho^{\mathcal{N}_{k}},G_{\mathcal{N}_{k}})$ be a copresheaf with the maps $\rho_{y\to x}^{\mathcal{N}_{k}}\colon\mathcal{F}^{\mathcal{N}_{k}}(y)\to\mathcal{F}^{\mathcal{N}_{k}}(x)$ define the transformations associated to the copresheaf. Given features $\mathbf{h}_{x}^{(\ell)}$, the next layer features are defined as
$\mathbf{h}_{x}^{(\ell+1)}=\beta\Biggl{(}\mathbf{h}_{x}^{(\ell)},\bigotimes_{k=1}^{n}\bigoplus_{y\in\mathcal{N}_{k}(x)}\alpha_{\mathcal{N}_{k}}\bigl{(}\mathbf{h}_{x}^{(\ell)},\rho_{y\to x}^{\mathcal{N}_{k}}(\mathbf{h}_{y}^{(\ell)})\bigr{)}\Biggr{)},$
where $\alpha_{\mathcal{N}_{k}}$ is the message function, $\oplus$ a permutation-invariant aggregator over neighbors $y\in\mathcal{N}_{k}(x)$, $\otimes$ combines information from different neighborhoods, and $\beta$ is the update function.
Proposition 3 lays the foundational framework that unifies a broad class of topological deep learning architectures, seamlessly bridging higher-order message passing methods and sheaf neural networks. This synthesis not only consolidates existing approaches but also opens avenues for novel architectures rooted in topological and categorical abstractions. Notably, the formulation in Proposition 3 encompasses simplicial message passing (Ebli et al., 2020; Bunch et al., 2020), cellular message passing (Hajij et al., 2020), stable message passing via Hodge theory (Hayhoe et al., 2022), and recurrent simplicial architectures for sequence prediction (Mitchell et al., 2024). It also subsumes more recent developments that leverage multiple signals and higher-order operators such as the Dirac operator (Calmon et al., 2022; Hajij et al., 2023a). These diverse models are unified under the copresheaf-based formulation by interpreting neighborhood aggregation, feature transport, and signal interaction within a coherent framework. See Appendix for derivations showing how several of these architectures emerge as special cases of this general formulation.
Remark 1 (Graph vs. CC copresheaf models).
Unlike graph-based models, which propagate information edge by edge, copresheaf models on a CC aggregate messages across all overlapping neighborhoods at once. Overlapping neighborhoods have common cells, potentially at different ranks, allowing simultaneous aggregation of multi-way interactions. Applying each neighborhood function $\mathcal{N}_{k}$ in turn, we compute its map-driven messages and then merge them into a single update.
4 Architectures Derived from the Copresheaf Framework
Having established the abstract copresheaf-based framework on a combinatorial complex $\mathcal{X}$, we now present several concrete instantiations. Copresheaf transformers (CTs) extend the standard attention mechanism by dynamically learning linear maps $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ encoding directional, anisotropic relationships between tokens, i.e., cells, in $\mathcal{X}$. Integrating these maps into attention enables CTs to capture rich, structured interactions. Copresheaf graph neural networks (CGNNs) generalize message passing GNNs by incorporating copresheaf linear maps to model relational structures. Copresheaf Convolutional Networks define convolution-like operations on combinatorial complexes, modeled as Euclidean grid, using these linear maps. We present the CT construction next and leave the exact formulations of copresheaf networks, CGNNs, and copresheaf convolution layers to the appendix.
Copresheaf transformers
Having established the abstract framework of CTNNs on a combinatorial complex $\mathcal{X}$, we now introduce a concrete instantiation: the Copresheaf Transformer (CT) layer. This layer extends the standard attention mechanism by dynamically learning linear maps $\rho_{y\to x}:\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{x}}$ that encode both the combinatorial structure and directional, anisotropic relationships within the complex. By integrating these maps into the attention computation, the CT layer captures rich, structured interactions across $\mathcal{X}$. At layer $\ell$, each cell $x\in\mathcal{X}$ is associated with a feature $\mathbf{h}_{x}^{(\ell)}\in\mathbb{R}^{d_{x}}$, where $d_{x}$ denotes the feature dimension of cell $x$.
Proposition 4 (Copresheaf self-attention).
For a fixed rank $k$ and neighborhood $\mathcal{N}_{k}$ (e.g., adjacency between $k$-cells), let $W_{q},W_{k}\in\mathbb{R}^{p\times d},\quad W_{v}\in\mathbb{R}^{d\times d}$ denote learnable projection matrices, where $p$ is the dimension of the query and key spaces, and $d$ is the feature dimension (assumed uniform across cells for simplicity). For each $k$-cell $x\in\mathcal{X}^{k}$, copresheaf self-attention defines the message aggregation and feature update as $\mathbf{h}_{x}^{(\ell+1)}=\beta\bigl{(}\mathbf{h}_{x}^{(\ell)},m_{x}\bigr{)}$, where $m_{x}=\sum_{y\in\mathcal{N}_{k}(x)}a_{xy}\rho_{y\to x}(v_{y})$ and
$a_{xy}=\frac{\exp(\langle q_{x},k_{y}\rangle/\sqrt{p})}{\sum_{y^{\prime}\in\mathcal{N}_{k}(x)}\exp(\langle q_{x},k_{y^{\prime}}\rangle/\sqrt{p})},$ (7)
where $q_{x}=W_{q}h_{x}^{(\ell)},k_{x}=W_{k}h_{x}^{(\ell)},~{}\mathrm{and}~{}v_{x}=W_{v}h_{x}^{(\ell)}$. Here, the the softmax normalizes over all neighbors $y^{\prime}\in\mathcal{N}_{k}(x)$ and $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ is the learned copresheaf map. The update function $\beta$ is chosen to be a neural network.
Similarly, we define copresheaf cross-attention among $s$ and $t$ rank cells in $\mathcal{X}$ detailed in our appendix, as well as a general algorithm for a corpresheaf transformer layer.
5 Experimental Evaluation
We conduct synthetic and real experiments in numerous settings to support the generality of our framework. These include learning physical dynamics, graph classification in homophillic and heterophilic cases and classifying higher-order complexes.
5.1 Evaluations on Physics Datasets
To verify the validity of our networks in toy setups of different characters, we generate a series of synthetic datasets. These include:
  1. 1.
    Heat. We generate 600 realisations by solving the heat equation $u_{t}=\nu u_{xx}$ on $[0,1)$ with $\nu=0.1$ to horizon $T=0.1$; each $u_{0}$ is a 10-mode sine series and $u_{T}$ is its Gaussian-kernel convolution, sampled on $N=100$ grid points.
  2. 2.
    Advection. Similar to heat, we generate 600 realisations of $u_{t}+c\,u_{x}=0$ with $c=1$; the solution is a pure phase shift of $u_{0}$, sampled on $N=130$ points. Pairs are min–max scaled to $[0,1]$ and split $500{:}100$ for train/test.
  3. 3.
    Unsteady stokes. Let $\mathbf{u}(x,y,t)=\bigl{(}u(x,y,t),\,v(x,y,t)\bigr{)}\in\mathbb{R}^{2}$ denote the incompressible velocity field, $p(x,y,t)\in\mathbb{R}$ the kinematic pressure, and $\nu>0$ the kinematic viscosity. Throughout, $\partial_{t}$ is the time derivative, $\nabla=(\partial_{x},\partial_{y})$ the spatial gradient, and $\Delta=\partial_{xx}+\partial_{yy}$ the Laplacian operator. The periodic unsteady Stokes system reads $\partial_{t}\mathbf{u}\;-\;\nu\,\Delta\mathbf{u}\;+\;\nabla p=\mathbf{0}$ where $\nabla\!\cdot\!\mathbf{u}=$. We synthesize $200$ samples by drawing an 8-mode Fourier stream-function $\psi$, setting $\mathbf{u}_{0}=\nabla^{\!\perp}\psi$, and evolving to $T=0.1$ with $\nu=0.1$ via the analytic heat kernel. Each pair is sampled on a grid $16\times 16$, channelwise min-max scaled to $[0,1]$, and split $160{:}40$ (train test).
Table 1: Mean squared error (mean $\pm$ std. dev.) of classical vs. copresheaf architectures for learning various physics simulations.
Network
Heat
Advection
Unsteady stokes
(Conv-transformer)
[…]
Model and training
For heat and advection, we leverage two transformer layers (positional encoding, 4 heads, stalk-dim $=16$) $\rightarrow$ followed by a mean pooling and linear head yielding 64D token embeddings. We train our networks using AdamW ($10^{-3}$) with a cosine LR scheduling, and batch size of two. We use 50 epochs for Heat dataset and 80 for Advection and report the results over three seeds. For Unsteady stokes, we test a compact Conv–Transformer U-Net: $2\!\rightarrow\!32$ conv encoder $\rightarrow$ two Transformer layers (4 heads, $d=32$, stalk-dim $=8$) $\rightarrow$ conv decoder back to two channels. We train it for 300 epochs using AdamW with a batch size of four. The classical baselines always use dot-product attention whereas the copresheaf variants employ learned outer-product maps.
Results
As shown in Tab. 1, in heat and advection tests copresheaf attention significantly outperforms classical dot-product attention, reducing the test MSE by over $50\%$ and achieving more stable results across seed. For unsteady stokes, the copresheaf attention lowers the test MSE by $\approx 15\,\%$ and reduces variance by an order of magnitude, confirming that pair-specific linear transports capture viscous diffusion of vorticity more faithfully than standard self-attention under identical compute budgets.
5.2 Graph Classification
We evaluate whether incorporating copresheaf structure into graph neural networks (GNNs) improves performance on graph classification tasks.
Data
MUTAG dataset, a nitroaromatic compound classification benchmark consisting of 188 molecular graphs, where nodes represent atoms and edges represent chemical bonds. Each node is associated with a 7-dimensional feature vector encoding atom type, and each graph is labeled as mutagenic or non-mutagenic (2 classes). The dataset is split into 80% training and 20% testing, with experiments repeated over multiple runs to ensure robustness.
Baselines, backbone and training
We compare standard GNN models (GCN, GraphSAGE, GIN) against their copresheaf-counterparts (CopresheafGCN, CopresheafSage, CopresheafGIN) derived below. All models are two-layer networks with a hidden dimension of 32 for GCN and GraphSAGE, and 16 for GIN, followed by global mean pooling and a linear classifier to predict graph labels. The standard models (GCN (Kipf and Welling, 2017b), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2019)) use conventional GNN convolutions: GCN with symmetric normalization, GraphSAGE with mean aggregation, and GIN with sum aggregation. The copresheaf-enhanced models (CopresheafGCN, CopresheafSage, CopresheafGIN) augment these with learned per-edge copresheaf maps, introducing local consistency constraints via transformations. All models are trained using Adam (lr=0.01) with a batch size of 16. GCN and GIN models are trained for 100 epochs, while GraphSAGE for 50. The negative log-likelihood loss is minimized, and performance is evaluated via test accuracy. For GCN and GraphSAGE, we use 5 runs, while GIN uses 10 runs.
Enhancing GNNs via copresheaves
The copresheaf structure enhances each GNN by learning a transport map $\rho_{ij}=I+\Delta_{ij}$ for each edge $(i,j)$, where $\Delta_{ij}$ is a learned transformation and $I$ is the identity matrix. The process for each model is as follows:
Table 2: Mean test accuracy ($\pm$ std).
Model Accuracy
GCN 0.674 $\pm$ 0.014
CopresheafGCN 0.721 $\pm$ 0.035
GraphSAGE 0.689 $\pm$ 0.022
CopresheafSage 0.732 $\pm$ 0.029
GIN 0.700 $\pm$ 0.039
CopresheafGIN 0.724 $\pm$ 0.021
  • CopresheafGCN. For node features $\mathbf{h}_{i},\mathbf{h}_{j}$, compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a diagonal $D\times D$ matrix. Form $\rho_{ij}=I+\Delta_{ij}$. Aggregate neighbor features as $\mathbf{h}_{i}^{\prime}=\sum_{j\in\mathcal{N}(i)}(\rho_{ij}\mathbf{h}_{j})/{\sqrt{d_{i}d_{j}}}$, where $d_{i},d_{j}$ are degrees. Combine with the self-feature: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$.
  • CopresheafSage. Compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a diagonal matrix, and form $\rho_{ij}=I+\Delta_{ij}$. Aggregate via mean: $\mathbf{h}_{i}^{\prime}=\text{mean}_{j\in\mathcal{N}(i)}(\rho_{ij}\mathbf{h}_{j})$. Combine: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$. The map $\rho_{ij}$ enhances local feature alignment.
  • CopresheafGIN. Compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a full $D\times D$ matrix, and form $\rho_{ij}=I+\Delta_{ij}$. Aggregate: $\mathbf{h}_{i}^{\prime}=\sum_{j\in\mathcal{N}(i)}(\rho_{ij}\mathbf{h}_{j})/d_{i}$. Combine: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$. The map $\rho_{ij}$ enables rich structural transformations.
Results
Copresheaf-enhanced GNNs consistently outperform their standard counterparts on the MUTAG dataset across all three architectures (GCN, GraphSAGE, GIN), with CopresheafSage showing the highest average accu- racy (0.732) and the largest relative improvement over GraphSAGE (0.689). The learned per-edge transport maps in copresheaf models enhance the ability to capture complex graph structures, improving classification performance by enforcing local consistency. These results highlight the potential of copresheaf structures in improving GNN performance for molecular graph classification tasks.
5.3 Combinatorial Complex Classification
Finally, we assess whether incorporating copresheaf structure into transformer-based attention mechanisms improves performance on classifying higher-order, general data structures, such as combinatorial complexes (CCs). Specifically, we compare a standard transformer model (Classical) against two CT variants (CT-FC and CT-SharedLoc) on a synthetic dataset of CCs.
Data
Our synthetic dataset comprises 200 training and 50 test CCs derived from Erdős-Rényi graphs, each with 10 nodes and a base edge probability of 0.5. Triangles are added to form higher-order structures with probability $q=0.1$ for class 0 (low density) or $q=0.5$ for class 1 (high density). Nodes have 2-dimensional feature vectors, consisting of their degree and the number of triangles it participates in, with added Gaussian noise ($\mathcal{N}(0,0.1)$).
Backbone and training
All models are transformer-based classifiers with a single block, using 2 attention heads, an embedding dimension of 8, and a head dimension of 4. The model embeds 2D node features, applies attention, performs global average pooling, and uses a linear classifier to predict one of two classes (low or high triangle density). The classical model employs standard multi-head attention. The CT-FC and CT-SharedLoc models augment attention with learned per-edge transport maps to enforce local consistency. Models are trained for 4 epochs using Adam (learning rate $10^{-3}$) with a batch size of 8, minimizing cross-entropy loss. Performance is evaluated via test accuracy. Experiments are over 4 runs with different random seeds to ensure robustness.
Copresheaf attention
Similar to GNNs, the copresheaf structure enhances attention by learning a transport map $\rho_{ij}=I+\Delta_{ij}$. The value vector $\mathbf{v}_{j}$ is transformed as $\rho_{ij}\mathbf{v}_{j}$ before attention-weighted aggregation. The process for each model is as follows:
  • CT-FC. For node features $\mathbf{h}_{i},\mathbf{h}_{j}$, compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a full $d\times d$ matrix ($d=4$). Form $\rho_{ij}=I+\Delta_{ij}$. Apply $\rho_{ij}$ to value vectors in attention: $\mathbf{v}_{t}=\rho_{ij}\mathbf{v}_{j}$. The map $\rho_{ij}$ enables rich feature transformations across nodes.
  • CT-SharedLoc. Compute a shared $\Delta_{ij}=\tanh(\text{MLP}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a full $d\times d$ matrix, and a local scalar $\alpha_{ij}=\sigma(\text{MLP}([\mathbf{h}_{i},\mathbf{h}_{j}]))$. Form $\rho_{ij}=I+\alpha_{ij}\Delta_{ij}$. Apply $\rho_{ij}$ to value vectors: $\mathbf{v}_{t}=\rho_{ij}\mathbf{v}_{j}$. The map $\rho_{ij}$ balances shared transformations with local modulation.
Results
Copresheaf transformer models outperform the standard transformer on the CC classification
Table 3: Mean and std test accuracy for CC classification.
Model Accuracy
Classic 0.940 $\pm$ 0.014
CT-FC 0.955 $\pm$ 0.009
CT-SharedLoc 0.970 $\pm$ 0.010
task, with CT-SharedLoc achieving the highest average accuracy (0.970) and competitive stability (std 0.010). The learned per-edge transport maps $\rho_{ij}$ enhance the model’s ability to capture higher-order structural patterns, such as triangle density, by aligning node features effectively. CT-SharedLoc’s combination of shared transport maps and local modulation yields the best performance, underscoring the value of copresheaf structures in transformer-based models for combinatorial complex classification.
6 Discussion and Conclusions
We proposed CTNNs, a unified deep learning framework on (un)structured data. By grounding models in copresheaves over CCs, CTNNs generalize GNNs, sheaf networks, and topological models through directional, heterogeneous message passing. Besides theoretical advances, CTNNs offer empirical benefits across diverse tasks laying the principles for multiscale and anisotropic representation learning.
Limitations and future work
CTNNs incur additional overhead from per-edge transformations and have so far been evaluated in modest-scale settings. We plan to explore well-engineered scalable parameterizations, extend CTNNs to large-scale and dynamic domains, and further connect categorical structure with robustness and inductive bias in deep learning.
Acknowledgments and Disclosure of Funding
Professor Adrian J. Lew’s contributions to this publication were as a paid consultant and were not part of his Stanford University duties or responsibilities. T. Birdal acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant number MR/Y018818/1] as well as a Royal Society Research Grant RG/R1/241402.
A Generalized Deep Learning Framework
–Supplementary Material–
Contents
[…]
Appendix A Notation
This appendix provides a reference summary of the notation and acronyms used throughout the main text. Table 4 details key mathematical symbols, while Table 5 lists abbreviations and their expansions.
Table 4: Summary of key notation used throughout the paper.
Notation Description $\mathcal{S}$ Underlying vertex set of a combinatorial complex (Def. 1) $\mathcal{X}$ Set of nonempty cells $\subseteq\mathcal{P}(\mathcal{S})$ (Def. 1) $\mathcal{P}(\mathcal{S})$ Power set of $\mathcal{S}$ (Def. 1) $\operatorname{rk}:\mathcal{X}\to\mathbb{Z}_{\geq 0}$ Rank function on cells, mapping to non-negative integers (Def. 1) $\mathbb{Z}_{\geq 0}$ Non-negative integers $\mathcal{X}^{k}$$\{x\in\mathcal{X}:\operatorname{rk}(x)=k\}$, the $k$-cells (Def. 1) $\dim\mathcal{X}$ Dimension of the complex, $\max_{x\in\mathcal{X}}\operatorname{rk}(x)$ (Sec. 2.1) $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ Neighborhood function, mapping cells to sets of neighbor cells (Def. 2) $\mathcal{N}_{\text{adj}}$ Adjacency neighborhood function (Def. 2) $\mathcal{N}_{\text{inc}}$ Incidence neighborhood function (Def. 2) $\mathcal{X}_{\mathcal{N}}$ Effective support of $\mathcal{N}$, $\{x\in\mathcal{X}\mid\mathcal{N}(x)\neq\emptyset\}$ (Sec. 3) $\mathbb{G}\in\{0,1\}^{m\times n}$ Binary neighborhood matrix (Def. 3) $\mathbb{G}^{\mathcal{N}}$ Copresheaf neighborhood matrix, entries $\rho_{z_{i}\to y_{j}}$ or 0 (Def. 7) $\rho_{y\to x}$ Copresheaf morphism $\mathcal{F}(y)\to\mathcal{F}(x)$ for edge $y\to x$ (Def. 4) $\mathbb{A}_{r,k}$ Copresheaf adjacency matrix between $r$-cells (Def. 8) $\mathbf{B}_{r,k}$ Copresheaf incidence matrix between ranks $r$ and $k$ (Def. 8) $C^{k}(\mathcal{X},\mathcal{F})$ $k$-cochain space, $\bigoplus_{x\in\mathcal{X}^{k}}\mathcal{F}(x)$ (Sec. 3) $\operatorname{Hom}(\mathcal{F}(i),\mathcal{F}(j))$ Space of linear maps from $\mathcal{F}(i)$ to $\mathcal{F}(j)$ (Def. 7) $\mathbf{h}_{x}^{(\ell)}$ Feature vector at cell $x$ in layer $\ell$ (Prop. 1) $\alpha$ Learnable message function (Prop. 1) $\beta$ Learnable update function (Prop. 1) $\oplus$ Permutation-invariant aggregator (Prop. 1, Prop. 3) $G=(V,E)$ Directed or undirected graph (Sec. 2.2) $G_{\mathcal{N}}=(\mathcal{X}_{\mathcal{N}},E_{\mathcal{N}})$ Directed graph induced by $\mathcal{N}$, edges $y\to x$ if $y\in\mathcal{N}(x)$ (Def. 6) $\mathcal{F}(x)$, $\mathcal{F}(e)$ Stalks: vector spaces at vertex $x$ or edge $e$ (Def. 4, Def. 5) $\mathcal{F}_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$ Restriction map in a cellular sheaf for $x\leq e$ (Def. 5) $\delta$ Coboundary map, measures local disagreement in cellular sheaf (Sec. 2.2) $\Delta_{\mathcal{F}}$ Sheaf Laplacian, $\Delta_{\mathcal{F}}=\delta^{T}\delta$ (Sec. 2.2) $\mathcal{Y}$, $\mathcal{Z}$ Collections of cells, used in neighborhood matrices (Def. 3, Def. 7) $W_{q}$, $W_{k}$, $W_{v}$ Learnable projection matrices for queries, keys, values in copresheaf self-attention (Prop. 4) $q_{x}$, $k_{x}$, $v_{x}$ Query, key, and value vectors for cell $x$ in copresheaf self-attention (Prop. 4) $a_{xy}$ Attention coefficient for cells $x$ and $y$ in copresheaf self-attention (Prop. 4)
Table 5: List of acronyms used throughout the paper.
Acronym Expansion CC Combinatorial Complex CTNN Copresheaf Topological Neural Network CMPNN Copresheaf Message-Passing Neural Network SNN Sheaf Neural Network GNN Graph Neural Network CNN Convolutional Neural Network CT Copresheaf Transformer CGNN Copresheaf Graph Neural Network GCN Graph Convolutional Network GraphSAGE Graph Sample and Aggregate GIN Graph Isomorphism Network CopresheafGCN Copresheaf Graph Convolutional Network CopresheafSage Copresheaf Graph Sample and Aggregate CopresheafGIN Copresheaf Graph Isomorphism Network NSD Neural Sheaf Diffusion SAN Sheaf Attention Network MLP Multi-Layer Perceptron GAT Graph Attention Network CAM Copresheaf Adjacency Matrix CIM Copresheaf Incidence Matrix CNM Copresheaf Neighborhood Matrix
Appendix B Sheaves and Copresheaves on Graphs: A Category Theoretical Look
This appendix provides a category-theoretic exposition of sheaves and copresheaves, emphasizing their definitions within the language of category theory. Additionally, we illustrate the construction of a copresheaf neighborhood matrix through an explicit combinatorial example, instantiating the concepts developed in the main text.
B.1 Copresheaves
Before diving into the technical definition, it’s helpful to think of a copresheaf as a way of assigning data that flows along the structure of a graph—like signals along neurons, or resources in a network. In categorical terms, this structure formalizes the idea of consistently associating elements of some category $\mathcal{C}$.
Definition 9 (Copresheaf on a directed graph).
Let $G=(V,E)$ be a directed graph, and let $\mathcal{C}$ be a category. A copresheaf on $G$ is a functor $\mathcal{F}:G\to\mathcal{C}$, where the graph $G$ is regarded as a category whose objects are the vertices $V$, and whose morphisms are the directed edges $(x\to y)\in E$. When $\mathcal{C}=\mathbf{Vect}_{\mathbb{R}}$, this structure corresponds to a quiver representation.
B.2 Cellular Sheaves
Definition 10 (Cellular sheaf on an undirected graph).
Let $G=(V,E)$ be an undirected graph. Define the incidence poset $(P,\leq)$, where $P=V\cup E$, and the order relation is given by $x\leq e$ whenever vertex $x\in V$ is incident to edge $e\in E$. A cellular sheaf on $G$ with values in a category $\mathcal{C}$ is a functor $\mathcal{F}:P\to\mathcal{C}$, which assigns:
  • to each vertex $x\in V$, an object $\mathcal{F}(x)\in\mathcal{C}$;
  • to each edge $e\in E$, an object $\mathcal{F}(e)\in\mathcal{C}$;
  • to each incidence relation $x\leq e$, a morphism $\rho_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$, called a restriction map,
such that the functoriality condition is satisfied on composable chains in the poset.
B.3 Comparison Between Copresheaves and Cellular Sheaves
Copresheaves provide a versatile and powerful framework for machine learning applications across diverse domains, particularly excelling in scenarios where directional data flow and hierarchical dependencies are paramount. Unlike cellular sheaves, which are defined over undirected graphs and enforce consistency through restriction maps $\rho_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$, copresheaves operate on directed graphs, assigning learnable linear maps $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ along edges. This enables anisotropic information propagation, making them ideal for tasks such as physical simulations—where data flows asymmetrically, as in fluid dynamics or heat transfer—or natural language processing, where sequential word dependencies dominate. More importantly, the vertex-centric design, assigning vector spaces $\mathcal{F}(x)$ solely to vertices, aligns seamlessly with message-passing architectures like Graph Neural Networks (GNNs) and Topological Neural Networks (TNNs), allowing these maps to be parameterized and optimized during training. Empirical evidence from our experiments highlights that copresheaf-based models outperform traditional architectures in capturing complex dynamics, demonstrating their superior ability to model spatially varying patterns and long-range dependencies in general applications. See Table 6 for a summary of the comparison between copresheaves and cellular sheaves.
Furthermore, copresheaves enhance machine learning models with a principled approach to regularization and expressiveness, broadening their suitability across heterogeneous domains. Standard neural network regularizers, such as $\ell_{2}$ decay or dropout, can be readily applied to copresheaf maps, with optional structural losses like path-consistency ensuring morphism compositionality. This adaptability stands in stark contrast to the rigid cohomological constraints of cellular sheaves, which enforce local-to-global agreement through terms like $\|\mathcal{F}_{x\leq e}(\mathbf{h}_{x})-\mathbf{h}_{e}\|^{2}$, limiting their flexibility in domains with asymmetric relationships. Copresheaves, by learning edge-wise maps, offer greater expressiveness for tasks involving non-Euclidean or multi-scale data, as evidenced by the superior performance of CopresheafConv layers on grid-based tasks. This makes them particularly effective for applications such as image segmentation, 3D mesh processing, or token-relation learning, where traditional methods like Convolutional Neural Networks (CNNs) struggle with directional or hierarchical structures. Our experiments further corroborate that copresheaf-augmented models consistently improve accuracy and detail recovery across diverse tasks, positioning them as a more suitable and generalizable tool for machine learning applications spanning Euclidean and non-Euclidean domains alike.
Table 6: Comparison between copresheaves and cellular sheaves.
Aspect Copresheaf Cellular sheaf [Hansen and Ghrist, 2019d]
Graph type Directed graph $G=(V,E)$ Undirected graph $G=(V,E)$
Assigned to vertices Vector space $\mathcal{F}(x)$ for each $x\in V$ Vector space $\mathcal{F}(x)$ for each $x\in V$
Assigned to edges Linear map $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ for each directed edge $x\to y\in E$ Vector space $\mathcal{F}(e)$ for each edge $e\in E$
Associated maps Pushforward: moves data forward along edges Restriction: pulls data back from vertices to edges
Map direction $\mathcal{F}(x)\to\mathcal{F}(y)$ (source to target) $\mathcal{F}(x)\to\mathcal{F}(e)$ (vertex to incident edge)
Interpretation Nodes have local features; edges transform and transmit them Edges represent shared contexts; vertex features are restricted into them
Goal / Objective Learn and compose edge-wise feature-space maps Enforce coherence across by gluing local data
Typical Regularization Standard NN regularizers ($\ell_{2}$ decay, spectral-norm, dropout, norm); optional structure losses (path-consistency, holonomy) Agreement between restricted vertex features and the edge-stalk, e.g., $\|\mathcal{F}_{x\leq e}(\mathbf{h}_{x})-\mathbf{h}_{e}\|^{2}$
Use in learning Embedding-level message passing, directional influence, anisotropic information flow Compatibility across shared structures, enforcing local consistency, cohomological constraints
B.4 Copresheaf Neighborhood Matrix Example
We define a copresheaf neighborhood matrix (CNM) for a combinatorial complex with an incidence neighborhood, guiding the reader through the setup, neighborhood, graph, copresheaf, and matrix.
Setup
Consider a combinatorial complex $\mathcal{X}=(\mathcal{S},\mathcal{X},\operatorname{rk})$ with $\mathcal{S}=\{a,b,c\}$, cells $\mathcal{X}=\{\{a\},\{b\},\{c\},\{a,b\},\{b,c\}\}$, and ranks $\operatorname{rk}(\{a\})=\operatorname{rk}(\{b\})=\operatorname{rk}(\{c\})=0$, $\operatorname{rk}(\{a,b\})=\operatorname{rk}(\{b,c\})=1$. Thus, $\mathcal{X}^{0}=\{\{a\},\{b\},\{c\}\}$, $\mathcal{X}^{1}=\{\{a,b\},\{b,c\}\}$. See Figure 3.
Figure 3: (a) A combinatorial complex $\mathcal{X}=(\mathcal{S},\mathcal{X},\operatorname{rk})$ with $\mathcal{S}=\{a,b,c\}$, cells $\mathcal{X}=\{\{a\},\{b\},\{c\},\{a,b\},\{b,c\}\}$. The figure also shows the induced directed graph $G_{\mathcal{N}_{\text{inc}}}=(\mathcal{V}_{\mathcal{N}},E_{\mathcal{N}})$ from the incidence neighborhood structure on the combinatorial complex $\mathcal{X}$. Each arrow $z\to y$ represents a directed edge from a 1-cell to a 0-cell where $y\subset z$, and is associated with a linear map $\rho_{z\to y}$ as part of the copresheaf. (b) The copresheaf neighborhood matrix (CNM) $\mathbb{G}^{\mathcal{N}_{\text{inc}}}$, where rows are indexed by 0-cells $\{a\},\{b\},\{c\}$ and columns by 1-cells $\{a,b\},\{b,c\}$. The matrix entries are linear maps $\rho_{z\to y}$ when $z\in\mathcal{N}_{\text{inc}}(y)$, and 0 otherwise. This matrix supports directional feature propagation from 1-cells to 0-cells.
Incidence neighborhood
The incidence neighborhood $\mathcal{N}_{\text{inc}}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ is:
$\mathcal{N}_{\text{inc}}(x)=\{y\in\mathcal{X}\mid x\subset y\}.$
For 0-cells: $\mathcal{N}_{\text{inc}}(\{a\})=\{\{a,b\}\}$, $\mathcal{N}_{\text{inc}}(\{b\})=\{\{a,b\},\{b,c\}\}$, $\mathcal{N}_{\text{inc}}(\{c\})=\{\{b,c\}\}$. For 1-cells: $\mathcal{N}_{\text{inc}}(\{a,b\})=\mathcal{N}_{\text{inc}}(\{b,c\})=\emptyset$. The effective support is $\mathcal{X}_{\mathcal{N}}=\{\{a\},\{b\},\{c\}\}$.
Induced graph
We induce a directed graph $G_{\mathcal{N}}=(V_{\mathcal{N}},E_{\mathcal{N}})$, with:
$V_{\mathcal{N}}=\mathcal{X}_{\mathcal{N}}\cup\bigcup_{x\in\mathcal{X}}\mathcal{N}_{\text{inc}}(x)=\{\{a\},\{b\},\{c\},\{a,b\},\{b,c\}\}=\mathcal{X},$
$E_{\mathcal{N}}=\{y\to x\mid x\in\mathcal{X}_{\mathcal{N}},y\in\mathcal{N}_{\text{inc}}(x)\}=\{\{a,b\}\to\{a\},\{a,b\}\to\{b\},\{b,c\}\to\{b\},\{b,c\}\to\{c\}\}.$
Copresheaf
The $\mathcal{N}_{\text{inc}}$-dependent copresheaf assigns $\mathcal{F}(x)=\mathbb{R}^{2}$ to each $x\in V_{\mathcal{N}}$, and linear maps $\rho_{y\to x}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ for $y\to x\in E_{\mathcal{N}}$:
$\rho_{\{a,b\}\to\{a\}}=\begin{bmatrix}1&0\\ 0&0.5\end{bmatrix},\quad\rho_{\{a,b\}\to\{b\}}=\rho_{\{b,c\}\to\{b\}}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\quad\rho_{\{b,c\}\to\{c\}}=\begin{bmatrix}1&0\\ 0&0.75\end{bmatrix}.$
Neighborhood matrix
The CNM $\mathbb{G}^{\mathcal{N}}$ for $\mathcal{Y}=\mathcal{X}^{0}$, $\mathcal{Z}=\mathcal{X}^{1}$ is:
$[\mathbb{G}^{\mathcal{N}}]_{i,j}=\begin{cases}\rho_{z_{j}\to y_{i}}&\text{if }z_{j}\in\mathcal{N}_{\text{inc}}(y_{i}),\\ 0&\text{otherwise}.\end{cases}$
For $\mathcal{Y}=\{\{a\},\{b\},\{c\}\}$, $\mathcal{Z}=\{\{a,b\},\{b,c\}\}$:
$\mathbb{G}^{\mathcal{N}}=\begin{bmatrix}\rho_{\{a,b\}\to\{a\}}&0\\ \rho_{\{a,b\}\to\{b\}}&\rho_{\{b,c\}\to\{b\}}\\ 0&\rho_{\{b,c\}\to\{c\}}\end{bmatrix}.$
This matrix facilitates message passing from 1-cells to 0-cells, e.g., bond-to-atom feature propagation.
More generally, Figure 4 shows an illustrative example of the general setup of copresheaf higher-order message passing on a CC with multiple neighborhood functions.
Figure 4: Illustration of copresheaf higher-order message passing. Left-hand Side: Shows a central cell $x$ (as a circle) with arrows pointing to boxes labeled $\mathcal{N}_{1}(x),\mathcal{N}_{2}(x),\ldots,\mathcal{N}_{k}(x)$, representing the collection of neighborhood functions $\mathfrak{N}=\{\mathcal{N}_{k}\}_{k=1}^{n}$. Right-hand Side: Depicts the message-passing process for the same cell $x$. Each neighborhood $\mathcal{N}_{k}(x)$ produces an aggregated message using $\bigoplus_{y\in\mathcal{N}_{k}(x)}\alpha_{\mathcal{N}_{k}}(\rho_{y\to x}(h_{y}^{(\ell)}))$. These messages are then combined using the inter-neighborhood function $\beta$, shown as a box, with an arrow updating $x$.
Appendix C Expressive Power of CTNNs
C.1 Universal Approximation of $\mathcal{N}$-Dependent Copresheaves
Here, we demonstrate that multilayer perceptrons (MLPs) can approximate arbitrary copresheaf morphisms induced by neighborhood functions. This result ensures that the proposed sheaf-based model is sufficiently expressive to capture complex data interactions.
Proposition 5 (Universal approximation of $\mathcal{N}$-dependent copresheaves).
Let $\mathcal{X}$ be a finite combinatorial complex and $\mathcal{N}$ a neighborhood function on $\mathcal{X}$. Suppose $G_{\mathcal{N}}\;\longrightarrow\;\mathbf{Vect}_{\mathbb{R}}$ is an $\mathcal{N}$-dependent copresheaf with stalks $\mathcal{F}^{\mathcal{N}}(x)=\mathbb{R}^{d}$ and morphisms $\rho^{\mathcal{N}}_{y\to x}$. Let a feature map
$h\colon\mathcal{X}\to\mathbb{R}^{d},\quad x\mapsto\mathbf{h}_{x}$
be given so that the $2d$-vectors $(\mathbf{h}_{y},\mathbf{h}_{x})$ are pairwise distinct for every directed edge $y\to x$. Define
$A=\bigl{\{}(\mathbf{h}_{y},\mathbf{h}_{x})\;\big{|}\;y\to x\bigr{\}}\;\subset\;\mathbb{R}^{2d},\quad g\colon A\to\mathbb{R}^{d\times d},\quad g(\mathbf{h}_{y},\mathbf{h}_{x})=\rho^{\mathcal{N}}_{y\to x}.$
Then for any $\varepsilon>0$ there exists a multilayer perceptron $\Phi\colon\mathbb{R}^{2d}\to\mathbb{R}^{d\times d}$ with sufficiently many hidden units such that $\bigl{\|}\Phi(\mathbf{h}_{y},\mathbf{h}_{x})-\rho^{\mathcal{N}}_{y\to x}\bigr{\|}<\varepsilon\,\text{for all }y\to x.$
Proof.
Since $A$ is finite and its elements are distinct, the assignment $g\colon A\to\mathbb{R}^{d\times d}$ is well-defined.
Enumerate $A=\{a_{i}\}_{i\in I}$, choose disjoint open neighborhoods $U_{i}\ni a_{i}$, and pick smooth “bump” functions
$\varphi_{i}\colon\mathbb{R}^{2d}\to[0,1],\quad\varphi_{i}(a_{i})=1,\quad\mathrm{supp}(\varphi_{i})\subset U_{i}.$
Then the sum
$f(a)\;=\;\sum_{i\in I}g(a_{i})\,\varphi_{i}(a)$
is a smooth map $f\colon\mathbb{R}^{2d}\to\mathbb{R}^{d\times d}$ satisfying $f|_{A}=g$.
Since $A$ is finite, choose a compact set $K\subseteq\mathbb{R}^{2d}$ containing $A$, ensuring the applicability of the Universal Approximation Theorem. By that theorem, for any $\varepsilon>0$ there is an MLP $\Phi$ such that
$\sup_{a\in K}\bigl{\|}\Phi(a)-f(a)\bigr{\|}<\varepsilon.$
In particular, for each $a_{i}\in A$ we have
$\bigl{\|}\Phi(a_{i})-g(a_{i})\bigr{\|}=\bigl{\|}\Phi(\mathbf{h}_{y},\mathbf{h}_{x})-\rho^{\mathcal{N}}_{y\to x}\bigr{\|}<\varepsilon,$
for every $y\to x$. This completes the proof. ∎
Appendix D Sheaf Neural Networks Are Copresheaf Message-Passing Neural Networks
In this appendix, we prove Theorem 1 and Proposition 2, which demonstrate that existing sheaf neural networks (SNNs), including sheaf diffusion networks, are special cases of the copresheaf message-passing neural network (CMPNN) framework (Definition 1). Moreover, we summarize how other sheaf-based neural architectures align with our unifying message-passing framework (see Table LABEL:tab:sheaf-homp-summary).
Proof of Theorem 1.
We construct the bidirected graph $G^{\prime}$ and define the copresheaf $\mathcal{G}$ on it, then demonstrate the equivalence of the message-passing operations.
First, construct the bidirected graph $G^{\prime}=(V,E^{\prime})$ from $G=(V,E)$ by replacing each undirected edge $\{u,v\}\in E$ with two directed edges $(u,v),(v,u)\in E^{\prime}$. This ensures that $G^{\prime}$ retains the connectivity of $G$ while introducing explicit directionality.
Next, define the copresheaf $\mathcal{G}:G^{\prime}\to\mathbf{Vect}_{\mathbb{R}}$ as follows:
  • For each vertex $x\in V$, assign $\mathcal{G}(x)=\mathcal{F}(x)$, where $\mathcal{F}(x)$ is the vector space associated to $x$ by the cellular sheaf $\mathcal{F}$.
  • For each directed edge $(v,u)\in E^{\prime}$, corresponding to the undirected edge $e=\{u,v\}\in E$, define the linear morphism $\rho_{v\to u}:\mathcal{G}(v)\to\mathcal{G}(u)$ by $\rho_{v\to u}=\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}$, where $\mathcal{F}_{v\triangleleft e}:\mathcal{F}(v)\to\mathcal{F}(e)$ and $\mathcal{F}_{u\triangleleft e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are the restriction maps of $\mathcal{F}$, and $\mathcal{F}_{u\triangleleft e}^{\top}:\mathcal{F}(e)\to\mathcal{F}(u)$ is the adjoint with respect to an inner product on $\mathcal{F}(e)$.
Now, consider the SNN message-passing mechanism along the edge $e=\{u,v\}$. For a feature vector $\mathbf{h}_{v}\in\mathcal{F}(v)$ at vertex $v$, the message transmitted to vertex $u$ is given by $\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}(\mathbf{h}_{v})$.
In the copresheaf $\mathcal{G}$ on $G^{\prime}$, the morphism associated with the directed edge $(v,u)$ is $\rho_{v\to u}=\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}$. Applying this morphism, the message-passing operation in the CMPNN yields $\rho_{v\to u}(\mathbf{h}_{v})=\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}(\mathbf{h}_{v})$, which is identical to the SNN message.
To ensure that $\mathcal{G}$ is a well-defined copresheaf, observe that it assigns vector spaces to vertices and linear maps to directed edges in a functorial manner. Specifically, for each directed edge $(v,u)\in E^{\prime}$, the map $\rho_{v\to u}$ is a composition of linear maps and thus linear. The identity and composition properties are satisfied implicitly through the consistency of the sheaf restriction maps.
Therefore, the SNN message passing, which operates via intermediate edge spaces in $\mathcal{F}$, is equivalently represented as direct vertex-to-vertex message passing in the copresheaf $\mathcal{G}$ on $G^{\prime}$. This completes the proof. ∎
Proof of Proposition 2.
Compute:
$(I_{n}\otimes W_{1})\mathbf{H}=[W_{1}\mathbf{h}_{v}]_{v\in V},$
$\bigl{(}\Delta_{\mathcal{F}}\otimes I\bigr{)}(I_{n}\otimes W_{1})\mathbf{H}W_{2}=\left[\sum_{u\in V}L_{F,v,u}W_{1}\mathbf{h}_{u}W_{2}\right]_{v\in V},$
$\mathbf{h}_{v}^{+}=\mathbf{h}_{v}-\sum_{u\in\mathcal{N}(v)\cup\{v\}}W_{2}L_{F,v,u}W_{1}\mathbf{h}_{u},$
since $L_{F,v,u}=0$ for $u\notin\mathcal{N}(v)\cup\{v\}$.
Interpret $G$ as a directed graph with edges $u\to v$ for $u\in\mathcal{N}(v)$ and $v\to v$. Define: - Message function: $\alpha(\mathbf{h}_{v},\rho_{u\to v}\mathbf{h}_{u})=W_{2}L_{F,v,u}W_{1}\mathbf{h}_{u}$, - Morphisms: $\rho_{u\to v}$ implicitly encoded via $L_{F,v,u}$, - Aggregator: $\oplus$=$\sum$ = ∑, - Update: $\beta(\mathbf{h}_{v},m)=\mathbf{h}_{v}-m$.
Thus:
$\mathbf{h}_{v}^{+}=\beta\left(\mathbf{h}_{v},\sum_{u\to v}\alpha(\mathbf{h}_{v},\rho_{u\to v}\mathbf{h}_{u})\right),$
matching Definition 1. ∎
Table LABEL:tab:sheaf-homp-summary provides a summary of sheaf neural networks realized in terms of Definition 1.
Table 7: Unified message passing formulations of various sheaf neural networks using HOMP notation. The restriction maps $\rho_{y\to x}$ may be linear, data-driven, or attentional depending on the model.